### v2

# mycolorRGB219, 48, 122

Abstract
We study phase separation between coexisting active and passive fluids in three-dimensions, using numerical simulation and experiments. Chaotic flows of the active phase drive giant interfacial deformations, causing the co-existing phases to interpenetrate and generate a continuously reconfiguring bicontinuous morphology which persists over the lifetime of the active fluid. Active bicontinuous structures are dominated by sheet-like interfaces, in marked difference from passive liquid-liquid phase separation which is controlled by saddle-like surfaces. Activity and surface tension control the length scale of the bicontinuous structure. These results demonstrate how active stresses suppress the coarsening of conventional phase separation, generating steady-state reconfigurable morphologies not accessible with conventional surface-modifying agents or through quenching of transient phase separated structures.
1 Introduction
Bicontinuous structures consist of two distinct interconnected and interpenetrating phases, where each phase spans the entire sample. By maximizing the interfacial area while retaining connectivity, bicontinuity underpins vital mechanical and chemical functionalities in diverse natural and synthetic systems. For example, in living matter, the interconnected endoplasmic reticulum that spans the cell interior performs protein synthesis, lipid production, and molecular transport  (?, ?). The bone trabecular lattice balances the mechanical strength of the mineralized struts with nutrient transport through vascularized marrow channels  (?, ?). In synthetic systems, spinodally decomposed alloys derive exceptional strength from their interpenetrating nanoscale phases, while block copolymer gyroids form porous electrodes that optimize transport in batteries (?, ?, ?, ?).
A homogeneous fluid mixture spontaneously demixes into two distinct phases  (?), a dynamical process that can generate a bicontinuous structure  (?, ?). In systems in thermal equilibrium, however, bicontinuous structure is inherently transient as the two phases coarsen over time reaching bulk phase separation  (?, ?). Thus, preserving bicontinuity requires kinetic arrest of the coarsening. Established strategies include suppressing the fluidity of one of the two phases. This can be achieved by replacing one of the two fluids with an elastomer network (?), by polymerizing one of the phase-separating fluids during coarsening (?, ?, ?), or by vitrifying it through rapid cooling (?). Alternatively, the interface itself can be stabilized by adsorbing a high density of colloids on the interface to form mechanically stable “bijels”  (?, ?, ?). Finally, block copolymers offer an alternate method for generating bicontinous structures through chemistry, but the scale of the resulting structure is controlled by the size of the constituent molecules  (?, ?, ?).
Figure 1: Active 3D phase separation yields a percolating network-like structure. (A-B) Morphology of active phase separation in computer simulations from the early state to the steady state. Red represents the active phase ($\phi>0.5$). Simulation parameters: activity $|\alpha|=1.30,$ surface tension $\gamma=1.0,$ and active fraction $\phi_{a}=0.40$. (C) Experimental morphology of the active phase (red) at steady state. (D) An $x$-$y$ cross-section of a phase-separating experimental sample. The white arrows indicate the active flows in the dextran phase (red), obtained from Particle Imaging Velocimetry. (E) MTs (cyan) partition into the dextran phase (red). (F) MTs in the dextran phase form bundles due to the depletion effect. Kinesin clusters (KSA) bind to MT bundles driving their extension. The experimental samples (C-E) contain 1.8% PEO and 1.8% dextran, $\phi_{a}=0.37$, and are imaged at time $t=1.5$ h. The KSA concentration is 183 nM (C) and 92 nM (D, E).
Here we describe a fundamentally different pathway for generating robust and tunable bicontinuous structures that uses active fluids and relies on mechanical, rather than chemical, control of soft interfaces. Active fluids are composed of energy-consuming molecular units that collectively generate autonomous chaotic flows  (?, ?, ?, ?, ?). Such flows and associated active stresses can power non-equilibrium dynamics and active assembly of passive colloids or fibers (?, ?). They can also drive the deformation and actuation of soft interfaces and membranes  (?, ?, ?, ?, ?, ?, ?, ?, ?, ?). Here we show how active flows can transform conventional coarsening dynamics into a continuously reconfiguring bicontinuous active network, which persists for the lifetime of the active fluid and is stable over a wide range of volume fractions. The bicontinuous structure is controlled by activity and surface tension. The interfacial curvature of the active bicontinuous network is dominated by sheet-like geometries, in contrast to passive phase-separating fluids where bicontinuous structures are dominated by minimal saddle-like surfaces  (?, ?, ?, ?, ?).
The three-dimensional geometry studied here is qualitatively different from previous work which explored the role of microtubule-based activity on phase separation of PEO/dextran mixtures. One class of studies explored active phase separation in very thin chambers where the sample was effectively two dimensional  (?, ?). Another study examined the effect of activity on the shape and stability of a bulk-separated interface in a three dimensional sample  (?). Here, in contrast, we explore the morphology of the late stage phase separated state and demonstrate the stability of bicontinuous structures.
2 Results
2.1 Theoretical model of active phase separation
We model an immiscible mixture of active and passive fluids in three dimensions within a continuum framework that couples Cahn-Hilliard phase separation to active nematodynamics. The concentration field $\phi(\mathbf{r})$ distinguishes the passive ($\phi=0$) and the active ($\phi=1$) phase, while the nematic tensor $\mathbf{Q}(\mathbf{r})$ captures the local orientational order within the active fluid. These fields are coupled through the incompressible Stokes flow $\mathbf{v}(\mathbf{r})$ and its spatial gradients (Methods 4.1)  (?, ?, ?).
The phase separation between active and passive fluid is governed by the equilibrium surface tension $\gamma$. The sample composition is controlled by the active fraction $\phi_{a}=(1/V)\int d\mathbf{r}\;\phi(\mathbf{r})$, where $V=\int d\mathbf{r}$ is the total volume. The active stress tensor that drives Stokes flow is
$\boldsymbol{\sigma}^{\text{active}}=\alpha\phi\mathbf{Q},$ (1)
where $\alpha<0$ is the extensile activity responsible for the turbulent-like dynamics of the active ($\phi=1$) phase. We choose a parameter regime in which the nematic tensor $\mathbf{Q}(\mathbf{r})$ relaxes to the isotropic state ($\mathbf{Q}=0$) in the absence of activity. Above a critical activity, active stresses destabilize the quiescent isotropic state and generate autonomous flows  (?, ?, ?, ?).
In passive systems, the well-mixed state is unstable over a range of parameters, leading to a spinodal instability that generates spatially modulated structures that coarsen over time. Such dynamics reduces the interfacial energy, ultimately resulting in bulk phase separation  (?, ?). Activity profoundly alters this process  (?). At high activity, turbulent-like flows stir the fluid and impede the thermodynamically driven phase separation  (?, ?). Intermediate activities do not suppress phase separation, but arrest coarsening and influence the morphology of the phase separated state. In two dimensions, moderate activity can generate percolating dynamical filamentary networks with a characteristic length scale controlled by the balance of activity and surface tension  (?).
We study the active-passive mixture at intermediate activities, using the numerical model. Starting with an initially uniform state, we observe the emergence of spatially modulated structures, which reach a length scale that plateaus over time (Fig. 1A, B, Movie S1). The steady-state structures have a labyrinth like appearance, consisting of junction points that are connected by thread-like liquid bridges. The network is not static. The threads connecting different junctions appear to continuously rupture while new ones are generated. Such microscopic dynamics generates a “living” continuously reconfiguring network-like structure.
2.2 Experimental realization of active phase separation
To experimentally study active-passive phase separation, we combined a previously described passive phase-separating polymer mixture of polyethylene oxide (PEO) and dextran with an active fluid (Methods 4.2)  (?, ?, ?). The activity is generated by microtubules (MTs) and streptavidin-bound clusters of kinesin molecular motors (KSA) (Fig. 1F). Both protein-based components partition into the dextran-rich phase (Fig. 1E). The kinesin clusters bind to adjacent MTs, transforming the chemical energy from adenosine triphosphate (ATP) hydrolysis into interfilament sliding and bundle extension, which in turn drives large-scale turbulent flows (Fig. 1D).
[…]
Figure 2: Bicontinuous morphology persists over a wide range of volume fractions. (A) Connected active (red) and passive phase (blue) in simulations, with $\phi_{a}=0.40$ (B) Connectedness factor $f_{c}$ of the active and passive phases as a function of the volume fraction of the active phase $\phi_{a}$. The range where $f_{c}\approx 1$ for both phases highlighted in pink indicates bicontinuity. Simulation parameters: $|\alpha|=0.70,\gamma=1.0$. (C) Morphology of the connected active (red) and passive phase (blue) in the experiment. $\phi_{a}=0.37$. (D) The connectedness fraction $f_{c}\approx 1$ for intermediate values of $\phi_{a}$, indicating bicontinuity. Experimental samples initially contain 183 nM KSA, 1.8% PEO, and 1.8% dextran.
Figure 3: Active width $w_{a}$ decreases with activity and increases with surface tension. (A-B) $w_{a}$ decreases with activity $|\alpha|$ in simulations and with KSA concentration in experiments. (C-D) $w_{a}$ increases with surface tension $\gamma$ in simulations and with capillary length in experiments. In (A-D) the color of the dots represents the volume fraction $\phi_{a}$ of the active phase (see color bar) and shows that $w_{a}$ is essentially independent of $\phi_{a}$. (E-G) 3D visualization of the dextran phase (red) in experiments with increasing KSA concentration, showing a corresponding decrease of $w_{a}$. (H-J) 3D visualization of the dextran phase (red) in experiments with increasing capillary length, showing a corresponding increase of $w_{a}$. The experimental samples (B, E-G) contain 1.8% PEO and 1.8% dextran. KSA concentrations are increased from 42 nM (E), 92 nM (F), to 183 nM (G). Samples (D, H-J) contain 183 nM KSA. PEO and dextran concentrations are both increased from 1.8% (H), 1.9% (I), to 2.0% (J).
2.3 Characterizing the bicontinuous morphology
Numerical simulations and experiments demonstrate that 3D active phase separation generates a steady-state dynamical structure. Active and passive phases form distinct yet intercalated system-spanning networks. Taken together, they generate a bicontinuous structure (Fig. 2 A,C). To analyze the bicontinuity in simulation, we threshold the 3D structure into passive and active regions, corresponding to $\phi<0.5$ and $\phi>0.5$, respectively. The connectedness fraction, $f_{c}$, of each phase is defined as the volume of the largest connected active/passive component divided by the total volume of that component  (?).
To determine the range of bicontinuity, we vary the fraction of the active phase $\phi_{a}$, while keeping the activity and surface tension fixed. Bicontinuous structures, defined as those where $f_{c}\approx 1$ for both phases, are observed from $0.3<\phi_{a}<0.6$ (Fig. 2 B) in simulation. For $\phi_{a}<0.3$, bicontinuity is broken as the active domains disintegrate into isolated droplets. For $\phi_{a}>0.6$, the passive domains are broken up by active flows and become disconnected. The range of stable bicontinuous structures is controlled by activity and surface tension (Fig. S5). For example, doubling the activity generates stable bicontinuous structures over the entire regime of phase separation $0.0<\phi_{a}<1.0$.
Motivated by these predictions, we have explored the phase space of stable bicontinuous networks in experiments. We use a phase recombination method to keep the polymer concentration (surface tension) and the motor protein concentration (activity) fixed within each phase, while the active fraction $\phi_{a}$ varies between 0.16-0.60. We (1) prepare an initial mixture with controlled concentrations of PEO, dextran and KSA, then (2) centrifuge the sample to generate bulk phase separation, and (3) recombine the two separated phases into a new mixture with chosen $\phi_{a}$ (Methods 4.2).
In experiments, the bicontinous structure for the particular is observed over the range of volume fractions $0.3<\phi_{a}<0.5$ (Fig. 2D). For active fractions greater than $0.5$, the passive phase disintegrates into droplets, while for fractions less than $0.3$ the active phase breaks up into droplets, significantly decreasing $f_{c}$. Notably, the range of stable bicontinuous structures is not centered around $\phi_{a}=0.5$: bicontinuity persists for a low volume fraction of the active phase $\phi_{a}\approx 0.3$. In comparison, it disappears as soon as the active phase reaches the majority $\phi_{a}\approx 0.5$. This asymmetry suggests that the active phase containing MTs remains connected at a very low volume fraction, while the passive PEO-rich/MT-devoid phase more easily breaks apart into the droplets. This asymmetry is more pronounced in experiments than simulations. Intriguingly, similar asymmetry was observed in the directional invagination and break-up of a bulk three-dimensional active interface, where it was attributed to the nematic elasticity of the active fluid  (?).
2.4 Controlling the morphology: active width
Simulations show that the morphology of the bicontinuous network is controlled by activity and surface tension. To characterize this structure, we measure the width $w_{a}$ of the active network, defined as the volume to surface ratio of the largest connected active domain. We find that $w_{a}$ increases with decreasing activity and with increasing equilibrium surface tension. Furthermore, $w_{a}$ is essentially independent of the volume fraction $\phi_{a}$ (Fig. 3A, C), suggesting that intrinsic active processes control the length scale of the bicontinuous morphology.
To experimentally test these predictions, we vary the KSA concentration, which is a proxy for activity, and the PEO/dextran concentration, which controls the surface tension. Due to a density difference, the dextran phase gradually sediments due to gravity over tens of minutes, causing a variation in $w_{a}$ (Fig. S3). To minimize this effect, we continuously invert the top and bottom of the chamber by slowly rotating the samples. The sample is then moved from the rotation stage to the microscope and imaged for a maximum duration of 5 minutes (Methods 4.2). The effective surface tension is estimated from measurements of the capillary length for different PEO/dextran concentrations (Fig. S1). The experimental trends follow the predictions of the numerical model: increasing surface tension increases $w_{a}$; increasing activity decreases $w_{a}$ (Fig. 3 A-D). Furthermore, $w_{a}$ is largely independent of the volume fraction.
The behavior of the width $w_{a}$ suggests that the morphology is controlled by two competing mechanisms: the stretching of the interface due to extensile active flows and its relaxation controlled by surface tension. Extensile active forces are known to align the nematic director with the interface  (?, ?, ?). This alignment then drives active shear flows that tangentially stretch the interface. The stretching is countered by surface tension that favors minimizing the interface. In general we expect activity to also modify the surface tension, as shown for instance in scalar models of active phase separation  (?). Here we assume that the main effect of activity is to build up local nematic order and align the liquid crystal with the interface, and use the bare (passive) surface tension in the scaling arguments below. In the isotropic liquid crystalline phase, the active stretching rate can be approximated as
$\tau^{-1}_{\text{active}}\sim\frac{S(|\alpha|-\alpha_{c})}{\eta}\sim\frac{(|\alpha|-\alpha_{c})^{3/2}}{\eta}\;,$ (2)
where $\eta$ is the fluid viscosity, $S$ is the scalar nematic order parameter and $\alpha_{c}$ is the threshold above which activity drives local nematic order (?, ?) (Methods 4.1). We approximate the scalar order parameter with the mean value generated by activity, $S\sim\sqrt{(|\alpha|-\alpha_{c})}$ (?, ?). In three-dimensions, the rate at which the interface relaxes due to surface tension is given by
$\tau^{-1}_{\text{diffusive}}\sim\frac{\gamma}{V}\;,$ (3)
where $V$ is the total volume of the active domain. We write this as $V=w_{a}^{{d_{f}}}\ell^{\left(3-{d_{f}}\right)},$ where $w_{a}$ is the width of the active region , $\ell$ is set by mass conservation, and $d_{f}<3$ is controlled by the active process   (?). Tube-like filamentary structures correspond to $d_{f}=2$ whereas sheet-like structures, with one short axis, correspond to $d_{f}=1$. At steady state, the active width is determined by balancing active stretching and interfacial relaxation, $\tau_{\text{active}}^{-1}\sim\tau_{\text{diffusive}}^{-1}$, with the result
$w_{a}\sim\left(\dfrac{\gamma}{(|\alpha|-\alpha_{c})^{3/2}}\right)^{1/d_{f}}=\left(\dfrac{|\alpha|-\alpha_{c}}{\gamma^{2/3}}\right)^{-3/2d_{f}}\;.$ (4)
Using simulations, we have measured the active width as a function of both activity and surface tension. The collapse of the data supports a scaling law $w_{a}\sim\left((|\alpha|-\alpha_{c})/\gamma^{2/3}\right)^{\nu},$ with $\nu\approx-1.36$, corresponding to $d_{f}=-3/(2\nu)\approx 1.1$ (Fig. 4). This scaling suggests that the bicontinuous morphology is dominated by sheet-like active domains, whose width is governed by the balance of active stretching and passive relaxation.
Figure 4: Scaling law for the width $w_{a}$ in simulations. Steady state width $w_{a}=\text{Volume}/\text{Surface Area}$ of the connected active domain in the bicontinuous region as a function of activity $|\alpha|$ and equilibrium surface tension $\gamma$, for a fixed active fraction $\phi_{a}=0.40$. The main figure shows that the data collapse when plotted versus $(|\alpha|-\alpha_{c})/\gamma^{2/3}$ (Eq. 4). The black dashed line shows the best fit (via linear regression) to $w_{a}\sim(|\alpha|-\alpha_{c})/\gamma^{2/3})^{\nu}$ for $\nu=-1.36\pm 0.04$. The inset shows the unscaled active width as a function of activity, for various values of $\gamma$. Fixed parameters: $K=24.0,\alpha_{c}=2\eta r/\lambda\Gamma=0.2$.
2.5 Characterizing curvature and morphology
The scaling argument described above suggests the dominance of sheet-like structures. To investigate further, we quantify the surface curvature in both experiment and simulation. We compute the values of the two principal curvatures $\kappa_{1}(\mathbf{s})$ and $\kappa_{2}(\mathbf{s})$, where $\mathbf{s}$ is a point on the interface $\mathcal{S}$. We define $\kappa_{1}(s)<\kappa_{2}(s)$ for each $s$ (Fig. 5A) (Methods 4.3). We visualize the interface structure with its normalized Gaussian curvature $w_{a}^{2}\kappa_{1}(\mathbf{s})\kappa_{2}(\mathbf{s})$ for $\mathbf{s}\in\mathcal{S},$ where $w_{a}$ is the width of the active phase, as defined in the previous section. In both experiments and simulations, the interface generated by active phase separation is dominated by regions of normalized Gaussian curvature close to zero (Fig. 5 B,C). We contrast this with the transient morphology obtained in passive phase separation from the Cahn-Hilliard equation, which is dominated by saddle-like surfaces with negative Gaussian curvature  (?, ?, ?) (Fig. 5D).
Figure 5: Comparing the interface morphology in passive and active bicontinuous structures. (A) The two principal curvatures $\kappa_{1}$ and $\kappa_{2}$ are calculated everywhere on the interface. (B-D) 3D visualization of the normalized Gaussian curvature $\kappa_{1}\kappa_{2}w_{a}^{2}$ of the interface from experiment (B), simulations of the active/passive mixture(C) and simulations of the passive Cahn-Hilliard equation (D). The color indicates the value of the Gaussian curvature. (E-H) Distribution $\rho(\kappa_{1},\kappa_{2})$ of normalized principal curvatures $\kappa_{1}w_{a}$ and $\kappa_{2}w_{a}$. Different values of $\kappa_{1}$ and $\kappa_{2}$ correspond to different geometries (E). $\rho$ peaks at small values of $\kappa_{1}$ and $\kappa_{2}$ in experiments (F) and simulations (G), indicating sheet-like interfaces. $\rho$ peaks at $\kappa_{1}\ll 0,\kappa_{2}\gg 0$ in simulations of the Cahn-Hilliard equation (H), indicating a saddle-like interface. The experimental sample (B,F) contains 92 nM KSA, 1.8% PEO, 1.8% dextran, and $\phi_{a}=0.37$. Simulation parameters (C,G): $|\alpha|=1.30,\gamma=1.0,$ and $\phi_{a}=0.40$.
From the principal curvature values, we compute the joint probability distributions $\rho(\kappa_{1},\kappa_{2})$. Different regions in the $\kappa_{1}$-$\kappa_{2}$ plane indicate different structures (Fig. 5 E). In both experiments and simulations, the joint distribution is peaked at small values of both $\kappa_{1}$ and $\kappa_{2},$ which corresponds to sheet-like domains (Fig. 5 F,G). The vertical and horizontal branches centered around zero-curvature corresponds to cylinder-like surfaces, which are associated with the sheet edges. In comparison, the probability distribution obtained from the Cahn-Hilliard equation is characterized by large positive-negative pairs of principal curvatures correspond to saddle-like surfaces (Fig. 5H). The curvature analysis supports the finding that the connected active domain is dominated by sheet-like structures.
3 Discussion
We showed how activity transforms the coarsening domains of passive three-dimensional phase separation into a continuously reconfiguring bicontinuous structure that maintains its morphology. The bicontinuous morphology is observed across a wide range of volume fractions, while its width is controlled by activity and surface tension. The active bicontinuous structure is composed of sheet-like domains, in contrast to saddle-like interfaces observed in passive phase separation  (?, ?, ?, ?, ?).
Extensive previous work examined the effect on activity on phase separation has been explored from multiple perspectives. In one regime, also studied here, active isotropic or nematic liquid crystals coexist with a passive fluid  (?, ?, ?, ?, ?, ?). In such systems, active stresses are present throughout the bulk of the active fluid. System-spanning active networks over a wide range of volume fractions have been studied in two-dimensional simulations  (?). Building on these findings, we use simulations and experiments to demonstrate that bicontinuity is a ubiquitous feature of three-dimensional active phase separation. An important ingredient required for the formation of bicontinuous structures is active anchoring  (?), where the active units align along the interfaces to generate autonomous flows that stretch the active phase into sheet-like domains. This elongation continuously connects separate droplets into a network and dynamically counters the collapsing effect of surface tension. Theoretical work also explored a different regime, described by scalar models of active/passive mixtures, where the mixture is described in terms of the conserved concentration field, coupled to flow  (?, ?, ?). In this regime, active stresses are driven by gradients of the concentration field that are localized at the interface. Whether bicontinuous structures can be obtained in three-dimensional scalar models, perhaps as a result of active renormalization of surface tension, remains, to our knowledge, an open question.
The structures in active phase separation are determined by the competition between active stress and surface tension. We have shown that this competition generates bicontinuous morphology and controls the width of the active sheets. The bicontinuity, however, is only observed over a wide but well-defined range of activities. On the one hand, when activity is very small compared to surface tension, active flows cannot elongate the droplets, but only induce small interfacial deformations. In this case, one phase irreversibly turns into droplets through a Plateau-Rayleigh instability  (?). On the other hand, for very high activities, the energy input overcomes the effect of the thermodynamic chemical potential, suppressing the phase separation itself and generating a uniform mixture (?, ?). Furthermore, bicontinuous structures are only observed in a finite range of active fraction $\phi_{a}$ (Fig. 2). At low fractions of the active fluid elongated droplets cannot merge. In this regime, the system transitions from droplets to a uniform phase with decreasing surface tension, without an intermediate bicontinuous state  (?). Finally, gravity causes a non-uniform distribution of the two phases, generating extreme densities $\phi_{a}$ which do not allow bicontinuity  (?). Our work provides specific techniques and parameters to overcome these difficulties and achieve a controllable bicontinuous fluid.
Bicontinuous structures also arise in passive phase separation  (?). However, without active forces, passive restorative interfacial tension drives the system to a state with a minimal surface area. The characteristic length of conventional coarsening dynamics grows with a power law $w\propto t^{\nu_{p}}$, where $\nu_{p}$ varies between 1/3 and 1  (?, ?, ?). Therefore, bicontinuous structures in passive systems are transient. In contrast, the active stresses maintain a dynamical, bicontinuous steady state throughout the entire active fluid lifetime.
Biological systems are adaptable and self-healing. Using continuous energy input, they maintain dynamical steady states, adapt to damage, and resist degenerative processes. Our work demonstrates how using continuous energy input can generate controllable and highly dynamical biomimetic structures and morphologies that are not easily achievable in equilibrium.
4 Materials and Methods
4.1 Continuum model and simulation details
We use a continuum model of phase separation in a mixture of an active liquid crystal and passive fluid. The model couples Cahn-Hilliard dynamics to the hydrodynamic of a nematic liquid crystal. It is formulated in terms of coupled equations for the phase fraction $\phi$ of active fluid, the nematic order parameter $\mathbf{Q}$ and fluid flow $\mathbf{v}$, given by
$\partial_{t}\phi+\mathbf{v}\cdot\boldsymbol{\nabla}\phi=M\nabla^{2}\mu,$ (5)
$\partial_{t}\mathbf{Q}+\mathbf{v}\cdot\boldsymbol{\nabla}\mathbf{Q}+[\boldsymbol{\Omega},\mathbf{Q}]=\lambda\mathbf{A}+\frac{1}{\Gamma}\mathbf{H},$
$0=\eta\nabla^{2}\mathbf{v}-\boldsymbol{\nabla}P+\boldsymbol{\nabla}\cdot\boldsymbol{\sigma},$
where $\phi(\mathbf{r})\in[0,1]$. The Stokes flow $\mathbf{v}(\mathbf{r})$ and its gradients given by the strain rate $A_{ij}=(\partial_{i}v_{j}+\partial_{j}v_{i})/2$ and the vorticity tensor $\Omega_{ij}=(\partial_{i}v_{j}-\partial_{j}v_{i})/2$ couple $\phi$ and $\mathbf{Q}$.
Phase separation between the active $(\phi=1)$ and the passive fluid $(\phi=0)$ is driven by the chemical potential $\mu=\delta F_{\phi}/\delta\phi,$ where $F_{\phi}=(3\gamma/\xi)\int d\mathbf{r}~\phi^{2}(\phi-1)^{2}+\xi^{2}(\nabla\phi)^{2}/2$. Here $\gamma$ is the equilibrium surface tension and $\xi$ is the width of the equilibrium interface between the two phases. We characterize the composition of the fluid mixture using the (conserved) active fraction $\phi_{a}=\int d\mathbf{r}\;\phi(\mathbf{r}).$
[…]
The nematic field is defined everywhere, but in the absence of activity it relaxes to $\mathbf{Q}=0$ since $r>0$. Due to the flow alignment and vorticity couplings, self-generated flows result in non-zero local nematic order within the active ($\phi=1$) phase, as we discuss below.
The flow is governed by the Stokes equation with incompressibility, i.e, $\boldsymbol{\nabla}\cdot\mathbf{v}=0$ enforced via the pressure $P$. The stress tensor $(\boldsymbol{\sigma}=\boldsymbol{\sigma}^{c}+\boldsymbol{\sigma}^{a})$ consists of the capillary stress ($\boldsymbol{\nabla}\cdot\boldsymbol{\sigma}^{c}=-\phi\boldsymbol{\nabla}\mu$), and the active stress, $\boldsymbol{\sigma^{a}}$, proportional to the phase field $\phi$ and the local nematic order $Q_{ij}$ i.e.
$\boldsymbol{\sigma}^{a}=\alpha\phi\mathbf{Q}\;,$ (6)
with $\alpha<0$ corresponding to extensile activity. Within the active phase ($\phi=1$) self generated flows arise above a critical activity $\alpha_{c}=2\eta r/\lambda\Gamma$. For $|\alpha|>\alpha_{c},$ the positive feedback of flow alignment due to the active flows generated by perturbations of the nematic field leads to active turbulence (?, ?, ?, ?).
Unless otherwise specified, the parameter values are reported in time units of $\Gamma/u$, length units of $\sqrt{M\Gamma}$ and energy density in units of $u$. Throughout this paper, we fix the flow alignment parameter $\lambda=1.0$, viscosity $\eta=1.0$, interface width $\xi=\sqrt{3/2}$ and free energy parameters r=0.1,w=0r=0.1,w=0.111We only consider large nematic order generated by active stresses, away from the isotropic to nematic transition, and hence can safely ignore the cubic term that appears in 3D. We verify in simulations that including a non-zero cubic energy contribution does not affect the stability of the bicontinuous phase. We consider different values of the nematic elasticity $K,$ and systematically vary the activity $|\alpha|,$ the equilibrium surface tension $\gamma,$ and the active volume fraction $\phi_{a}.$ The simulations are done with periodic boundary conditions and performed using self-developed pseudo spectral solvers (?) on a uniform three-dimensional grid of size $170\times 170\times 170,$ typically for $10^{5}$-$10^{6}$ time steps until at least the system reaches steady state.
4.2 Experimental Design
The experimental system of active-passive phase separation is based on previous protocols  (?, ?), with a decreased concentration of coexisting polymers to reduce surface tension. PEO (100 kDa, Sigma-Aldrich) and dextran (450 to 650 kDa, Sigma-Aldrich) were added into M2B buffer (80 mM K-pipes, 2 mM MgCl2, 1 mM EGTA, pH 6.8). We studied three initial concentrations: 1.8% PEO and 1.8% dextran, 1.9% PEO and 1.9% dextran and 2.0% PEO and 2.0% dextran. Decreasing polymer concentration decreased the capillary length and surface tension (Fig. S1). For visualization, we added (<<0.1% w/w) of amino-dextran (Fina Biosolutions) fluorescently labeled with Alexa-Fluor 488 NHS Ester (Invitrogen).
Tubulin was purified and labeled as described previously  (?). A minor tubulin fraction (3%) was conjugated to AZDye 647 NHS Ester (Fluoroprobes) using a succinimidyl ester linkage. Labeled and unlabeled tubulin monomers were co-polymerized into microtubules (MTs) at 37 C using 0.6 mM guanylyl-(alpha,beta)-methylene-diphosphonate (GMPCPP, Gena Biosciences) and 1.2 mM Dithiothreitol (DTT, Gena Biosciences) in M2B buffer (?). The MTs were sonicated for 60 seconds in a Branson 1800 sonicator to reduce filament length, which generated softer bundles that fragment more easily. The final MT concentration was 0.67 mg/mL. Kinesin-401 protein fused to a biotin-carboxyl carrier domain was expressed and purified using established protocols  (?). Kinesin-401 was combined with tetrameric streptavidin in an M2B buffer with 5 mM DTT, forming a kinesin-streptavidin cluster (KSA). The final concentrations of KSA in the samples varied from 42-183 nM.
[…]
The denser dextran phase sediments in the z-z direction over tens of minutes. This would generate two bulk phase separated phases, with the active phase perforated with passive droplets at the bottom of the chamber. To minimize this effect, samples were placed on a rotator (Benchmark Scientific H2020 Roto-Therm) and rotated at 24 rpm at room temperature for 1.5 hours. This reduced sedimentation (Fig. S4). For imaging, the samples were transferred to the microscope and imaged for a maximum of 5 minutes (Fig. S3).
4.3 Data Analysis
The experimental data were analyzed using established methods (?). Confocal microscopy images were first classified into the dextran-rich phase and PEO-rich phase with the machine-learning software Ilastik based on intensity thresholding (?). The interface was extracted between the two phases to calculate the width of the structure (Fig. 3). The interface was then transformed into a triangular mesh form using the software MeshLab (?). The curvature was then calculated using published codes in MATLAB (?, ?). 3D visualization of both simulation and experiment data was done with software Imaris (Fig. 1,2,3) (?).
Supplementary materials
Supplementary Text
Figs. S1 to S6
Movies S1 to S2
Supplementary Materials for
Bicontinuity in active phase separation
Paarth Gulati1†, Liang Zhao1†, Michio Tateno2, Omar A. Saleh1,2,3,
Zvonimir Dogic1,4, M. Cristina Marchetti1,4∗ Corresponding author: mcmarche@physics.ucsb.edu
These authors contributed equally to this work.
This PDF file includes:
Supplementary Text
Figs. S1 to S6
Captions for Movies S1 to S2
Other Supplementary Materials for this manuscript:
Movies S1 to S2
Supplementary Text
Capillary length measurement
The capillary lengths $l_{c}$ between the PEO and dextran phase were measured using a previously published method (?). Passive samples containing no MT or KSA and containing 1.81.8%-2.02.0% PEO and 1.81.8%-2.02.0% dextran were placed in transparent fluorinated ethylene propylene (FEP) tubes with inner diameters $2.4$ mm. The samples were centrifuged at $2000$ g relative centrifugal force for $1$ min (Fisher Scientific, 05-090-128) to bulk separate the PEO-rich and dextran-rich phases. The samples were then imaged on a laser sheet microscope (Zeiss Z.1 Lightsheet). The dextran phases de-wetted the FEP wall (Fig. S1 A). The interface profiles near the wall were then fitted by exponential functions to obtain $l_{c}$ (Fig. S1 B).
[…]
Gravity effect on bicontinuous structure
In the experiment, since the dextran phase has a slightly larger density than the PEO phase, it gradually sediments in the z-z direction due to gravity over time. This causes an inhomogeneous amount of dextran phase at different $z$ positions (Fig. S3 B, C). To minimize this effect, we define the area fraction of the dextran phase at each $z$ as
$\phi_{2D}(z)=A_{d}(z)/A_{total}(z)$ (S1)
where $A_{d}(z)$ is the area of dextran phase and $A_{total}(z)$ is the area of the whole sample. In each sample, we pick the $z$ position where $\phi_{a}-0.05<\phi_{2D}(z)<\phi_{a}+0.05$ (Fig. S3 A) for all analysis.
To verify that the bicontinuous structure persists in steady state, we image the sample at tt=10 min, 40 min, 120 min and 180 min (Fig. S4). In each acquisition, we image the sample for 5 min, and we keep the sample on a rotor at 24 rpm between the acquisitions. We observe the bicontinuity persists and $w_{a}$ remains the same over 2 hours (Fig. S4), which reveals that the sample is in steady state in absence of sedimentation, consistent with simulation results (Fig. 1 B).
Figure S2: Passive phase separated samples form droplets (A,B) Two dimensional crosssections of passive samples with initially 1.81.8% PEO and 1.81.8% dextran, no KSA. The passive phase forms droplets when $\phi_{a}=0.56>0.5$ in (A) and the active phase form droplets when $\phi_{a}=0.41<0.5$ in (B).
Figure S3: Gravity causes inhomogeneous dextran phase fraction at different $z$ positions. (A) Dextran area fraction $\phi_{2D}$ vs. $z$, averaged over 5 minutes. $z=0$ is the middle of the sample. The gravitational field points along negative $z$. The volume fraction of dextran phase $\phi_{a}=0.48$. The black box indicates where $\phi_{a}-0.05<\phi_{2D}(z)<\phi_{a}+0.05$. (B, C) 2D horizontal cross-sections of the sample in (A) at $t=2.5$ min and $z$ position where $\phi_{2D}(z)=\phi_{a}+0.05$ (B) and $\phi_{2D}(z)=\phi_{a}-0.05$ (C). The sample has an initial concentration of $42$ nM KSA, 1.81.8% PEO and $1.8$ % dextran.
Figure S4: Bicontinuous structure reached steady state. 3D visualizations of the dextran phase (red) at 4 different time points (A-D). The network structure reached a steady state at $t=40$ min and remained similar over 2 hours (B-D) together with the active width $w_{a}$ (E). The sample initially contain 92 nM KSA, 1.8% PEO and 1.8% dextran and is recombined to obtain active fraction $\phi_{a}=0.28$.
Figure S5: Bicontinuity for increasing activity. The dependence of active/passive domain connectedness fractions on the activity. For large activities, the bicontinuous state for the entire range of volume fractions explored numerically i.e. the spinodal region of the passively phase separating system.
Figure S6: Visualization of active phase (red) morphology over varying $\phi_{a}$. Samples contain $\phi_{a}=$0.16 (A), 0.29 (B), 0.37 (C), 0.60 (D), corresponding to data points in Fig. 2D. All samples initially contain 183 nM KSA, 1.8% PEO, and 1.8% dextran and are recombined to obtain various $\phi_{a}$.
Caption for Movie S1.
Three-dimensional visualization of active phases (red) forming dynamic networks in experiment at different KSA concentrations. The movie starts at $1.5$ h after all sample were prepared. The samples initially contain 1.81.8% PEO, 1.81.8% dextran, 4218342-183 nM KSA.
Caption for Movie S2.
Three-dimensional visualization of active phase (red) forming a dynamic network in simulation. The sample is same as Fig. 1 A-B.